# Self-improving estimates of growth of subharmonic and analytic functions

### v1

@@ 2. Our main results
then there exist positive constants $c_{1},c_{2}$ and $\tau$ such that
@@
Let $0<{p_{*}}<k$ and let $\Omega$ be a bounded open subset of $\mathbb{R}^{k}$. Suppose that $B\subset\mathbb{R}^{k}$ is closed and $A\subset\Omega$ is a compact $p_{*}$}-admissible set such that $m(A)=0$. Suppose that $g(1/t)$ belongs to a Hardy field $\mathfrak{R}$, which contains all functions $t^{\gamma}$, $\gamma\in\mathbb{R}$. Suppose also that for some $\beta>0$ and $C_{1}>0$, $g(t)\leq C_{1}t^{-\beta}$ for $t\in(0,1]$. Then for any $a>1$ there exists a constant $v\in(0,1)$ such that the following holds. For any subharmonic function $u$ on $\Omega\setminus B$ satisfying (1.1), one has
@@
    Let $g(t)=t^{-b}$, $0<b<q$. Then, by Theorem 2.5, (2.2) implies that $u(x)\leq Cg(\operatorname{dist}(x,B))$, $x\in\Omega\setminus B$. If $A$ is a Lipschitz curve, $k\geq 3$ and $b\leq k-2$, one can apply Theorem 2.3 to get the same conclusion.
@@
    Still consider the case $g(t)=\log^{b}(1/t)$, where $0<a<\infty$ and $A$ is a Lipschitz curve.
@@
$A\cap\mathcal{D}_{a,R}=\mathcal{G}(\varphi)=\big{\{}(x_{1},x^{\prime\prime}):\;x^{\prime\prime}=\varphi(x^{\prime}),a^{\prime}-R<x^{\prime}<a^{\prime}+R\big{\}}$
@@ 4. Quantitative versions of Domar’s theorems
$\eta^{-}(s):=\inf\{t\in(\alpha,\beta):\;\eta(t)\leq s\},\quad s\in(\textstyle{\lim_{\beta}\eta},\infty).$
@@ 4.1. Our version of Domar’s Theorem A
Let us see that Lemma 4.3 can in fact be used to obtain a quantitative version of Theorem A. First, notice that the function $\delta\colon(\lambda+1,\infty)\to\mathbb{R}$ is positive, strictly decreasing and $\delta(t)\to\infty$ when $t\to(\lambda+1)^{+}$. On the other hand, the function $\psi\colon(\lambda+1,\infty)\to\mathbb{R}$ is decreasing. Moreover, if
@@
We will use the right-continuous function $\varphi^{-}\colon(0,\infty)\to(\lambda+1,\infty)$.
@@
Suppose that (4.2) and (4.6) hold and define the function $\varphi\colon(\lambda+1,\infty)\to\mathbb{R}$ as in Lemma 4.3. Then for all $x\in\Omega$ we have
@@ 4.2. Our version of Domar’s Theorem B
Suppose that $\mu_{q_{*}}(\nu)$ is finite for any $\nu>0$ and (4.9) holds. Define the function $\rho\colon(\lambda,\infty)\to\mathbb{R}$ as in Lemma 4.7. Then for all $x\in\Omega$ we have
@@ 1. $(\mathcal{A}(p,C))$
$u(x)\leq a^{\rho_{ad}^{-}(\operatorname{dist}(x,B)/(3D))}\leq ag\big{(}v\operatorname{dist}(x,B)\big{)},\quad\operatorname{dist}(x,B)\leq\tau$ (5.8)
@@
Can the conclusion (2.4) of Theorem  2.5 be substituted by an estimate

### v2

@@ 2. Our main results
then there exist positive constants $c_{1},c_{2}$ and $\tau$, which do not depend on the function $u$, such that
@@
Let $0<{p_{*}}<k$ and let $\Omega$ be a bounded open subset of $\mathbb{R}^{k}$. DY Suppose that $B\subset\mathbb{R}^{k}$ is closed and $A\subset\Omega$ is a compact $p_{*}$}-admissible set such that $m(A)=0$. Let $g:(0,+\infty)\to(0,+\infty)$ be a strictly decreasing function such that $g(0^{+})=+\infty$ and $g(t)\leq C_{1}t^{-\beta}$ for $t\in(0,1]$, where $\beta>0$ and $C_{1}>0$ are constants. Suppose also that $g(1/t)$ belongs to a Hardy field $\mathfrak{R}$, which contains all functions $t^{\gamma}$, $\gamma\in\mathbb{R}$. Then for any $a>1$ there exists a constant $v\in(0,1)$ such that the following holds. For any subharmonic function $u$ on $\Omega\setminus B$ satisfying (1.1), one has
@@
    Let $g(t)=t^{-b}$, $0<b<q$. Then, by Theorem 2.5, (2.2) implies that $u(x)\leq Cg(\operatorname{dist}(x,B))$, $x\in\Omega\setminus B$. If $A$ is a closed subset of a Lipschitz curve, $k\geq 3$ and $b\leq k-2$, one can apply Theorem 2.3 to get the same conclusion.
@@
    Still consider the case $g(t)=\log^{b}(1/t)$, where $0<b<\infty$ and $A$ is closed subset of a Lipschitz curve.
@@
$A\cap\mathcal{D}_{a,R}=\mathcal{G}(\varphi)=\big\{(x_{1},x^{\prime\prime}):\;x^{\prime\prime}=\varphi(x^{\prime}),a^{\prime}-R<x^{\prime}<a^{\prime}+R\big\}.$
@@ 4. Quantitative versions of Domar’s theorems
$\eta^{-}(s):=\inf\{t\in(\alpha,\beta):\;\eta(t)\leq s\},\quad s\in(\textstyle{\lim_{\beta}\eta},\infty).$ (4.1)
@@ 4.1. Our version of Domar’s Theorem A
Let us see how Lemma 4.3 can be used to obtain a quantitative version of Theorem A. First, notice that the function $\delta\colon(\lambda+1,\infty)\to\mathbb{R}$ is positive, strictly decreasing and $\delta(t)\to\infty$ when $t\to(\lambda+1)^{+}$. On the other hand, the function $\psi\colon(\lambda+1,\infty)\to\mathbb{R}$ is decreasing. Moreover, if
@@
We will use the right-continuous function $\varphi^{-}\colon(0,\infty)\to(\lambda+1,\infty)$ (see (4.1)).
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{k}$, $k\geq 2$, and let $F\colon\Omega\to[0,\infty]$ be a nonnegative upper semicontinuous function such that $m(\{x\in\Omega:F(x)>s\})>0$ for all $s$, and
$\int_{\Omega}\big[\log^{+}F(x)\big]^{k-1+\varepsilon}\,dx<\infty$
for some $\varepsilon>0$. Let $\mathcal{C}_{F}$ be the class of all subharmonic functions $u\colon\Omega\to\mathbb{R}$ such that $u\leq F$ and set $M(x):=\sup\{u(x):u\in\mathcal{C}_{F}\}$. Fix a constant $a>1$ and let the function $\varphi$ be as in Lemma 4.3. Then for all $x\in\Omega$ we have
@@ 4.2. Our version of Domar’s Theorem B
Let $\Omega$ be a bounded domain in $\mathbb{R}^{k}$, $k\geq 2$, and let $F\colon\Omega\to[0,\infty]$ be a nonnegative upper semicontinuous function. Fix a constant $a>1$ and let $F_{\nu}:=\{x\in\Omega:F(x)\geq a^{\nu}\}$ for each $\nu>0$. Choose some ${p_{*}}\in(0,k)$ and put ${q_{*}}=k-{p_{*}}$. Set
$\mu_{q_{*}}(\nu):=\sup\,\frac{m(F_{\nu}\cap B_{k}(x,R))}{R^{p_{*}}},\qquad\nu>0,$
where the supremum is taken over all balls $B_{k}(x,R)$ contained in $\Omega$. Suppose that $\mu_{q_{*}}(\nu)$ is finite for any $\nu>0$ and
$\int_{0}^{1}\mu_{q_{*}}^{-}(s)\,ds^{1/{q_{*}}}<\infty$
(see (4.1)). Define the function $\rho$ as in Lemma 4.7. Then for all $x\in\Omega$ we have
@@ 1. $(\mathcal{A}(p,C))$
$u(x)\leq a^{\rho_{ad}^{-}(\operatorname{dist}(x,B)/(3D))}\leq ag\big(v\operatorname{dist}(x,B)\big),\quad\operatorname{dist}(x,B)\leq\tau,$ (5.8)
@@
Can the conclusion (2.4) of Theorem 2.5 be substituted by an estimate